# references.bib

### v1

@@ Abstract
Recently, a class of mechanical lattices with reconfigurable, zero-stiffness structures has been proposed, called Totimorphic structures. In this work, we introduce a computational framework that allows continuous reprogramming of a Totimorphic lattice’s effective properties, such as mechanical and optical properties, via continuous geometric changes alone. Our approach is differentiable and guarantees valid Totimorphic lattice configurations throughout the optimisation process, thus providing not only specific configurations with desired properties but also trajectories through configuration space connecting them. It enables re-programmable structures where actuators are controlled via automatic differentiation on an objective-dependent cost function, altering the lattice structure at all times to achieve a given objective – which is interchangeable to achieve different functionalities. Our main interest lies in deep space applications where harsh, extreme, and resource-constrained environments demand solutions that offer flexibility, resource efficiency, and autonomy. We illustrate our framework through two proofs of concept: a re-programmable metamaterial as well as a space telescope mirror with adjustable focal length, both made from Totimorphic structures. The introduced framework is easily adjustable to a variety of Totimorphic designs and objectives, providing a light-weight model for endowing physical prototypes of Totimorphic structures with autonomous self-configuration and self-repair capabilities.
Keywords morphing structure  $\cdot$ deployable structure  $\cdot$ multi-functional metamaterial  $\cdot$ totimorphic  $\cdot$ programmable material  $\cdot$ inverse design  $\cdot$ automatic differentiation  $\cdot$ in-orbit deployment  $\cdot$ space infrastructure  $\cdot$ telescope
@@ 1 Introduction
Throughout nature, the intricate and disordered lattice structures that are observed in bones, plant stems, dragonfly wings, coral, radiolarians [gibson2010cellular], amongst many other examples, demonstrate how powerful geometry is for designing structures with extreme mechanical properties from a very limited selection of base materials [wegst2015bioinspired]. Metamaterials [surjadi2019mechanical] are a recent example of human-engineered lattice structures that utilise the geometric design space of unit cells to change the properties of the lattice obtained by tiling this motive, often producing structures with different properties than those of the underlying lattice material – for instance, having a soft and compressible lattice made of a very brittle material such as ceramic [meza2014strong]. In addition to metamaterials that follow a periodic design philosophy, there is a growing interest in (inversely) designing disordered lattice materials and structures [aranguren2018designing, torres2019bone, ross2021using, maurizi2022inverse, dold2023differentiable, shumilin2023method, zheng2023unifying, jadhav2024generative], allowing us to fully tap into the functional design space explored by nature. Since lattices can be constructed using additive manufacturing, they combine ease of manufacturing with a highly expressive design space that only requires a small amount of building materials. It is not surprising that lattices have found applications on a variety of scales, ranging from nano- and mesoscale materials to large-scale structures such as space habitats [ghidini2018materials, park2023soll, gregg2024ultralight, thomas37totimorphic]. The static nature of lattices also means that once they have been constructed, their properties are fixed – unless physically stimulating the lattice changes the properties of its base materials or allows switching between different shapes (e.g., magnetically [montgomery2021magneto, wu2022magnetically, wang2023physics]), therefore enabling a certain degree of reprogrammability of the lattice’s properties; also known as active metamaterials [xiao2020active, xia2022responsive].
Recently, a type of reconfigureable lattice called a Totimorphic structure has been proposed and demonstrated in a physical prototype [chaudhary2021totimorphic]. The Totimorphic property of these lattice structures enables the continuous alteration of effective properties of the lattice through mechanically induced continuous geometric changes alone. Totimorphic lattices are made of a triangle-shaped unit cell composed of a beam (Fig. 1.A, points $\mathrm{A}$ to $\mathrm{B}$), a lever that connects via a ball joint to the middle of the beam (Fig. 1.A, points $\mathrm{P}$ to $\mathrm{C}$), and two zero-length springs that connect the two ends of the beam with the other end of the lever (Fig. 1.A, $\mathrm{A}$ to $\mathrm{C}$ and $\mathrm{B}$ to $\mathrm{C}$). Due to the zero-length property of the springs, the Totimorphic unit cell implements a mechanical version of Thales’s theorem: in any given lever position (lying on a half-circle), the forces from the springs pulling at the lever combine such that the total force is parallel to the lever orientation, resulting in no torque on the lever. Consequently, the unit cell in this state is neutrally stable: it is completely compliant to reconfiguration but will not move in the absence of external forces.
@@
Figure 1: Illustration of Totimorphic lattices. (A) Totimorphic unit cell and elementary cell used for the tile, as defined in this work. (B) Actuation of angles (or lattice dynamics) are defined as gradient descent on an objective-dependent cost function. This way, both a lattice configuration with a desired property is found as well as the trajectory connecting it to the initial configuration. The surface plot represents the cost for different configurations. An example trajectory following the gradient $-\nabla C$ is shown in white, and the corresponding lattice configurations to the right. For illustration purposes, the task here is to move the upper right corner of the lattice (blue dot) to the height of the red line. In more realistic scenarios, the landscape might contain insurmountable barriers, because the lattice needs to break in order to move between certain configurations.
Mathematically, the Totimorphic unit cell is characterised by a set of three geometric constraints, which we call Totimorphic condition in the remainder of the paper. The three constraints are as follows:
@@
    The beam has a fixed length $l_{\text{b}}$.
@@
    The lever has a fixed length $\frac{1}{2}l_{\text{b}}$.
@@
    The lever is connected at one end to the midpoint of the beam.
A Totimorphic unit cell is only allowed to transform through configurations that satisfy these constraints. When assembling unit cells into a lattice structure, it can morph into different shapes by adjusting beam and lever orientations as long as each unit cell satisfies the Totimorphic condition and unit cells stay connected, i.e., the lattice does not break.
In previous work [chaudhary2021totimorphic], the lattice is usually initialised as a flat surface. A desired surface shape is then found by optimising the lattice point positions using an objective-dependent, quadratic cost function, with the Totimorphic condition being accounted for as non-linear constraints. However, this comes with two limitations: firstly, the Totimorphic condition is never satisfied exactly, but only up to a certain error. Secondly, this approach only provides the final configuration with the desired shape, not the trajectory through configuration space to get there – or more simply put, we do not know how to morph from the initial to the found configuration without breaking the lattice. In the introduced approach, we avoid these problems by formulating the reconfiguration of Totimorphic lattices as an unconstrained optimisation problem. This is achieved by choosing a new set of parameters $\mathcal{P}$ to describe the lattice, from which the physical coordinates $\mathcal{R}$ (beam and lever ends, i.e., points $\mathrm{A}$, $\mathrm{B}$, $\mathrm{P}$, $\mathrm{C}$, and primed variants) of the lattice are calculated analytically, $\mathcal{R}=f_{\mathrm{T}}(\mathcal{P})$. These new parameters are chosen such that the map $f_{\mathrm{T}}$ from new to physical lattice coordinates always satisfies the Totimorphic condition. We derive new parametrisations and suitable maps both for lattices in two and three-dimensional space. Although this can be easily applied to any Totimorphic lattice [^1] In fact, the calculations required to map from generalized coordinates to lattice coordinates are always the same, so it is possible to implement a general Totimorphic model supporting arbitrary unit cell connectivity and tilings., we illustrate it for a specific tiling used in [chaudhary2021totimorphic] where the lattice is formed from hour-glass shaped cells (made from two Totimorphic unit cells), see Fig. 1.A. We call these cells elementary cells in the remainder to distinguish them from unit cells.
The utilized approach is known in analytical mechanics as generalized coordinates, which is often used to describe mechanical systems with constraints, or as the substitution method in mathematical optimisation. In our case, the chosen generalized coordinates are all real-valued and, different from the physical lattice coordinates, bounded. Additionally, we find that the map $f_{\mathrm{T}}$ is differentiable, and hence automatic differentiation (i.e. gradient descent) can be used to optimize cost functions written in terms of lattice and generalized coordinates. Consequently, performing gradient descent on the cost function is equivalent to following a trajectory through configuration space that morphs the lattice into a configuration exhibiting a desired mechanical or optical property without breaking it (Fig. 1.B); which can be efficiently and robustly realised using modern automatic differentiation libraries such as pyTorch. In other terms: when interpreting a Totimorphic lattice as a dynamic system that reconfigures continuously, its dynamics (or actuations) are given by the gradient of the cost function. Thus, our approach provides not only a target lattice configuration with desired properties, but also the trajectory connecting the initial lattice configuration with the target one.
We use the introduced framework to explore new application areas for Totimorphic lattices. Different from previous work which only focused on shape shifting [chaudhary2021totimorphic], we optimise the topology of Totimorphic lattices to continuously change its effective material properties. Specifically, we propose Totimorphic structures as deployable and reconfigurable meso - and macro-scale space infrastructure, where the focus lies on adapting the functionality of the structure through autonomous reconfiguration.
Figure 2: Totimorphic structures as building blocks of, e.g., space habitats, spacecraft components, and tools with reprogrammable properties. In this work, we specifically look into lattice structures that can alter their mechanical properties, i.e., how they behave under stress. In addition, we explore the concept of using Totimorphic structures as the backbone of large-scale space infrastructures such as space telescopes – not only to ease their deployment, but to equip them with reprogrammable properties as well. For example, we demonstrate a mirror element that can adjust its surface shape to alter the location of its focal point.
Several designs of reprogrammable materials have recently been proposed and verified in the aerospace field, with related examples being Mechanical Neural Networks [lee2022mechanical] – lattices with programmable stiffness of each beam elements – and the Trimorph origami pattern [liu2022triclinic] – an origami structure with triclinic symmetry that can transition between multiple discrete stable states with different mechanical properties such as Poisson’s ratios with opposite signs. As a first use case, we demonstrate Totimorphic lattices with continuous mechanical properties (Fig. 2, left), meaning that effective properties such as Poisson’s ratio or stiffness can be adjusted continuously through continuous changes in the configuration of the Totimorphic lattice, i.e., lever and beam orientations obtained via gradient descent.
As a second use case, we apply our framework to large scale space structures and spacecraft such as space telescopes, antennas, and habitats. A variety of futuristic concepts have recently been proposed to reduce the costs and risks of launching and deploying large space infrastructure. Particularly prominent ones are the use of formation flight [mettler2005large, ayre2005self], CubeSat swarms that assemble and reconfigure by docking into static formations, e.g., to form space telescope mirrors [pirat2022toward], autonomously assembling tiles for in-orbit habitat construction [ayre2005self, ekblaw2021self], Origami techniques to unfold spacecraft components [trease2013accommodating, zirbel2014deployment], as well as robotic systems for ground-based habitat construction on other celestial bodies [gregg2024ultralight]. Since a major constraint for the possible size and structure of spacecraft is the launcher used for deployment in orbit or deep space, such concepts are in action already, e.g., the James Webb Space Telescope was folded together for launch and subsequently unfolded [^2] https://webb.nasa.gov/content/about/launch.html##webbLaunchConfiguration, visited July 2024.. However, most concepts focus only on deployment, restricting (or at least severely limiting) any further reconfiguration of the spacecraft after deployment to adapt to changing mission parameters or changes in the environment the spacecraft is operating in. Here, we propose that Totimorphic lattices represent an intriguing candidate to form the backbone of deployable and reconfigurable large-scale space infrastructure. As a proof of concept, we use our framework to construct a Totimorphic mirror with reconfigurable focal length as well as self-repair capabilities (Fig. 2, right).
In summary, we introduce a computational framework for autonomously reconfiguring Totimorphic structures for different functional tasks, focusing on adapting the effective properties of lattices by changing only their geometric structure. Reconfiguration is realized as gradient descent minimizing an objective-dependent cost function, therefore enabling continuous movement through the configuration space of Totimorphic lattices to adjust their properties. We further highlight and validate in simulation the applicability of this framework in the space domain, adding to the growing number of concepts of self-constructing and reconfigurable designs crucial for future uncrewed deep space exploration.
2 Results
In the following, we first introduce how the generalised coordinates of a Totimorphic lattice are chosen and how they are used to analytically calculate the physical coordinates of the lattice elements. After introducing the model, we present the results of two proofs of concept: a meso-scale lattice material with continuously programmable material properties, as well as a large-scale space telescope with continuously adjustable optical properties. Mathematical details on all models can be found in Section 4. Simulation details are in the supplemental information.
Figure 3: Parametrizing Totimorphic lattices. (A) When connecting elementary cells to an existing lattice, overlapping elements are shared with already existing cells (shown in red). Thus, the number of degrees of freedom of the newly added cell is reduced. (B) Iterative construction of a lattice made of Totimorphic cells. (C) If all points except one in an elementary cell are predetermined (shown in blue), the remaining point is obtained from the Totimorphic condition. This is illustrated here for the case where $\mathrm{A^{\prime}}$ and $\mathrm{C}$ are given. Since $\mathrm{P^{\prime}}$ is connected via a beam and lever to these points, it has to be a distance of $\frac{1}{2}l_{\text{b}}$ away from both. The solution to this problem is given by the intersection points of two circles, or equivalently, by a quadratic equation. Two points $\mathrm{P^{\prime}}$ and $\mathrm{P}^{\prime\prime}$ satisfy the Totimorphic condition, however, the latter would reverse the order of $\mathrm{A^{\prime}}$ and $\mathrm{B^{\prime}}$ and is therefore not a valid physical solution.
2.1 Degrees of freedom in Totimorphic lattices
To model Totimorphic lattices, we absorb the Totimorphic condition into the parameters used to describe the lattice’s configuration. The minimum number of parameters needed to uniquely capture the lattice is given by its degrees of freedom $df$. In general, the degrees of freedom of an elementary cell in a Totimorphic lattice are determined by three factors: the freely moving points characterising the cell itself, the Totimorphic condition, and points that overlap with other cells in the lattice – with the latter two factors reducing the number of degrees of freedom. In the following, we construct a lattice cell by cell to identify how a cell’s degrees of freedom depend on where and when it is added to the lattice. For simplicity, we showcase this for lattices in two dimensions:
  1. 1.
    The first elementary cell that we add is not constrained by any neighbouring cells (Fig. 3.A1). It is characterised by five points: $\mathrm{A}_{0}$, $\mathrm{B}_{0}$, $\mathrm{C}_{0}$, $\mathrm{A}^{\prime}_{0}$, and $\mathrm{B}^{\prime}_{0}$. In two dimensions, this yields ten parameters (two for each point). However, due to the Totimorphic condition, i.e. the constant lengths of the four beams, there are only six degrees of freedom (each length constraint reduces the degrees of freedom by one), and so we only require six parameters to fully describe the elementary cell.
  2. 2.
    Adding another elementary cell above the first one, the points $\mathrm{A}_{1}$ and $\mathrm{B}_{1}$ are already determined, $\mathrm{A}_{1}=\mathrm{A}^{\prime}_{0}$ and $\mathrm{B}_{1}=\mathrm{B}^{\prime}_{0}$ (Fig. 3.A2). Thus, we have six parameters (point $\mathrm{C}_{1}$, $\mathrm{A}^{\prime}_{1}$ and $\mathrm{B}^{\prime}_{1}$), which leaves us with three degrees of freedom for this elementary cell when accounting for the three remaining length constraints (two levers and one beam). The same is true when adding an elementary cell below the first one.
  3. 3.
@@ 2.1 Degrees of freedom in Totimorphic lattices
    Adding another elementary cell to the right of the first one, both points $\mathrm{A}_{2}$ and $\mathrm{A}^{\prime}_{2}$ are already determined, $\mathrm{A}_{2}=\mathrm{B}_{0}$ and $\mathrm{A}^{\prime}_{2}=\mathrm{B}^{\prime}_{0}$ (Fig. 3.A3). Since all four length constraints (two levers and two beams) still have to be enforced, we are left with only two degrees of freedom in this case. The same is true when adding an elementary cell to the left of the first one.
@@
    Adding an elementary cell above and to the right of two other cells, the points $\mathrm{A}_{3}$, $\mathrm{B}_{3}$ and $\mathrm{A}^{\prime}_{3}$ are already determined, $\mathrm{A}_{3}=\mathrm{B}^{\prime}_{0}$, $\mathrm{B}_{3}=\mathrm{B}^{\prime}_{1}$, $\mathrm{A}^{\prime}_{3}=\mathrm{B}^{\prime}_{1}$ (Fig. 3.A4). Since three length constraints still have to be satisfied, we are left with only one degree of freedom (four from the remaining points minus the three length constraints). This is in general true when adding an elementary cell such that two of its adjacent sides are connected to other elementary cells.

### v2

# Continuous Design and Reprogramming of Totimorphic Structures for Space Applications

@@ Abstract
Recently, a class of mechanical lattices with reconfigurable, zero-stiffness structures has been proposed, called Totimorphic lattices. In this work, we introduce a computational framework that enables continuous reprogramming of a Totimorphic lattice’s effective properties, such as mechanical and optical behaviour, through geometric changes alone, demonstrated using computer simulations. Our approach is differentiable and guarantees valid Totimorphic configurations throughout the optimisation process, providing not only target states with desired properties but also continuous trajectories in configuration space that connect them. This enables reprogrammable structures in which actuators are controlled via automatic differentiation on an objective-dependent cost function, continuously adapting the lattice to achieve a given goal. We focus on deep space applications, where harsh and resource-constrained environments demand solutions that combine flexibility, efficiency, and autonomy. As proof of concept, we present two scenarios: a reprogrammable disordered lattice material and a space telescope mirror with adjustable focal length. The introduced framework is adaptable to a wide range of Totimorphic designs and objectives, providing a lightweight model for endowing physical systems with autonomous self-configuration and self-repair capabilities.
Keywords morphing structures $\cdot$ deployable structures $\cdot$ Totimorphic lattices $\cdot$ programmable material $\cdot$ inverse design $\cdot$ automatic differentiation $\cdot$ in-orbit deployment $\cdot$ space infrastructure $\cdot$ space telescopes
@@ 1 Introduction
Figure 1: Illustration of Totimorphic lattices reconfigured using automatic differentiation. (A) Totimorphic unit cell. For tiling, we use the shown elementary cell as the basic building block in this work. (B) Actuation of generalized coordinates (here: beam and lever angles) is defined as gradient descent on an objective-dependent cost function $C$. This way, both a lattice configuration with a desired property is found as well as the trajectory connecting it to the initial configuration. The surface plot represents the cost for different configurations. An example trajectory following the gradient $\nabla C$ is shown in white (from 1 to 4), and the corresponding lattice configurations to the right. For illustration purposes, the task here is to move the upper right corner of the lattice (blue dot) to the height of the red line. In realistic scenarios, the landscape might contain holes, i.e., configurations that break the structure.
Self-assembling and reconfigurable space infrastructure has the potential to reduce deployment costs and risks while increasing mission flexibility and long-term reusability. A variety of concepts for in-orbit assembly have lately been proposed, including formation flight [15, 1], CubeSat swarms that dock into space mirrors [18], autonomously assembling tiles for habitat construction [1, 5], origami-based deployment of spacecraft components [29, 36], and robotic systems for ground-based assembly on other celestial bodies [9]. However, most of these approaches focus solely on deployment and offer little or no capability for post-deployment reconfiguration, and are therefore unable to adapt to new tasks, repair damage, or respond to shifting environmental conditions autonomously. Recent advances such as Mechanical Neural Networks [12] and Trimorph origami patterns [13] begin to explore programmable materials with reprogrammable properties, but remain limited to discrete configurations or narrow ranges of tunable properties.
Nature offers powerful examples of how disordered geometry alone can endow structures with exceptional mechanical properties [31]. Prominent examples are bones, plant stems, dragonfly wings, coral, and radiolarians [7], and the concept is further found in human-engineered metamaterials [25, 16, 33, 32]. Inspired by recent work on inversely designing irregular mechanical lattices [34, 28, 22, 14, 4, 24, 35, 11], we propose a new concept for reconfigurable space infrastructure with a high degree of programmability: continuously inverse-designed Totimorphic lattices.
Totimorphic lattices are a recently introduced type of reconfigureable mechanical lattice [3]. Their triangular unit cell consists of a beam (fig.˜1A, points $\mathrm{A}$ to $\mathrm{B}$), a lever attached at the beam’s midpoint (points $\mathrm{P}$ to $\mathrm{C}$), and two zero-length springs (springs where the spring force is proportional to the total length of the spring) that connect the lever tip to the beam ends (wiggly lines). The beam-lever connection and connections between unit cells use ball (pin) joints. Because of this arrangement, the spring forces always balance without exerting torque on the lever. As a result, each unit cell — and any lattice assembled from them — is neutrally stable: it can reconfigure freely under external actuation, yet remains stationary when unloaded.
In previous work [3], target lattice shapes were obtained through constrained optimisation, with beam and lever lengths enforced as constraints. This approach has two main drawbacks: constraints are only satisfied approximately, and just the final configuration is produced, not the trajectory through configuration space to get there. In [27], we demonstrated shape-morphing, e.g., from a flat surface sheet into a half-cylinder, using a marching cell method, which however struggled with maintaining lattice integrity and could become stuck while morphing between shapes. Here we address both issues by introducing generalised coordinates that absorb the geometric constraints into the parameterisation itself. The resulting mapping from generalised to physical lattice coordinates is differentiable, allowing us to use automatic differentiation on task-specific cost functions. This way, Totimorphic lattices are reconfigured continuously along valid trajectories until the cost function is minimised (fig.˜1B).
We validate our framework in the space domain using simulations, adding to the growing number of concepts of self-constructing and reconfigurable designs. As a first demonstration, we show that Totimorphic lattices can continuously tune their effective mechanical properties, making them feasible as building blocks for adaptive structures with tunable stiffness or Poisson’s ratio (Fig. 2, left). As a second, we explore their use in large-scale infrastructure by simulating a reconfigurable space mirror capable of adapting its focal length and repairing structural damage (Fig. 2, right).
In the following, we first introduce the generalised coordinates and differentiable reconfiguration framework, followed by the two proof of concept space applications. Mathematical details on all models, proofs, and simulation details are found in section˜4. Simulation code is available online on GitHub [8].
Figure 2: Totimorphic lattices as building blocks of, e.g., space habitats, spacecraft components, and tools with reprogrammable properties. In this work, we specifically look into lattice structures that can alter their mechanical properties, i.e., how they behave under stress. In addition, we explore the concept of using Totimorphic lattices as the backbone of large-scale space infrastructures such as space telescopes – not only to ease their deployment, but to equip them with reprogrammable properties as well. For example, we demonstrate a mirror element that can adjust its surface shape to alter the location of its focal point.
2 Results
2.1 Degrees of freedom in Totimorphic lattices
To model Totimorphic lattices, we absorb the beam and lever length / angle constraints into the parameters used to describe the lattice’s configuration. The minimum number of parameters needed to uniquely capture the lattice is given by its degrees of freedom $df$. In general, the degrees of freedom of an elementary cell in a Totimorphic lattice are determined by three factors: the freely moving points characterising the cell itself, the beam/lever constraints, and points that overlap with other cells in the lattice – with the latter two factors reducing the number of degrees of freedom. In the following, we construct a lattice cell by cell to identify how a cell’s degrees of freedom depend on where and when it is added to the lattice. For simplicity, we showcase this for lattices in two dimensions:
@@ 2.1 Degrees of freedom in Totimorphic lattices
    The first elementary cell that we add is not constrained by any neighbouring cells (Fig. 3A1). It is characterised by five points: $\mathrm{A}_{0}$, $\mathrm{B}_{0}$, $\mathrm{C}_{0}$, $\mathrm{A}^{\prime}_{0}$, and $\mathrm{B}^{\prime}_{0}$. In two dimensions, this yields ten parameters (two for each point). However, due to the beam/lever constraints, i.e. the constant lengths of the four beams, there are only six degrees of freedom (each length constraint reduces the degrees of freedom by one), and so we only require six parameters to fully describe the elementary cell.
@@
    Adding another elementary cell above the first one, the points $\mathrm{A}_{1}$ and $\mathrm{B}_{1}$ are already determined, $\mathrm{A}_{1}=\mathrm{A}^{\prime}_{0}$ and $\mathrm{B}_{1}=\mathrm{B}^{\prime}_{0}$ (Fig. 3A2). Thus, we have six parameters (point $\mathrm{C}_{1}$, $\mathrm{A}^{\prime}_{1}$ and $\mathrm{B}^{\prime}_{1}$), which leaves us with three degrees of freedom for this elementary cell when accounting for the three remaining length constraints (two levers and one beam). The same is true when adding an elementary cell below the first one.
@@
    Adding another elementary cell to the right of the first one, both points $\mathrm{A}_{2}$ and $\mathrm{A}^{\prime}_{2}$ are already determined, $\mathrm{A}_{2}=\mathrm{B}_{0}$ and $\mathrm{A}^{\prime}_{2}=\mathrm{B}^{\prime}_{0}$ (Fig. 3A3). Since all four length constraints (two levers and two beams) still have to be enforced, we are left with only two degrees of freedom in this case. The same is true when adding an elementary cell to the left of the first one.
@@
    Adding an elementary cell above and to the right of two other cells, the points $\mathrm{A}_{3}$, $\mathrm{B}_{3}$ and $\mathrm{A}^{\prime}_{3}$ are already determined, $\mathrm{A}_{3}=\mathrm{B}^{\prime}_{0}$, $\mathrm{B}_{3}=\mathrm{B}^{\prime}_{2}$, $\mathrm{A}^{\prime}_{3}=\mathrm{B}^{\prime}_{1}$ (Fig. 3A4). Since three length constraints still have to be satisfied, we are left with only one degree of freedom (four from the remaining points minus the three length constraints). This is in general true when adding an elementary cell such that two of its adjacent sides are connected to other elementary cells.